OpenStudy (anonymous):

A helicopter is trying to drop supplies in the light green drop zone. If it lands in the dark green area, the supplies will get tangled in trees. What is the probability that the supplies will land in the trees? (Assume that it won't land outside the colored shape.)

OpenStudy (anonymous):



OpenStudy (anonymous):

biggest one is light green

OpenStudy (anonymous):

i need help setting it up

OpenStudy (danjs):

area of the tree zone : total area

OpenStudy (anonymous):

i got 3/8?

OpenStudy (danjs):

which one is the trees?

OpenStudy (danjs):

3 / 11

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

whats 3/11??

OpenStudy (danjs):

(area of trees) / (total area)

OpenStudy (anonymous):

im confused

OpenStudy (danjs):

The area of the bigger box is 4*2 = 8 The area of the small box is 3*1 = 3 The total area is the sum of them = 8 + 3 = 11

OpenStudy (danjs):

So the probablility of hitting the trees is the area of the trees (3) out of the entire area of the trees and the drop zone (11)

OpenStudy (anonymous):

11/3 or 3/11 so probable